OpenStudy (lxelle):

The parametric equations of a curve are x = 2theta + sin 2 theta, y=1-cos 2 theta. Show that dy/dx = tan theta.

OpenStudy (irishboy123):

dy / dx = {dy/dt} / {dx/dt}

OpenStudy (lxelle):

dt is dtheta?

OpenStudy (irishboy123):

yep

OpenStudy (lxelle):

dx/dt would be 2-2cos2theta?

OpenStudy (irishboy123):

if x = theta + sin 2 theta as per question stated above then: dx/dtheta = 1 + 2 cos 2 theta

OpenStudy (lxelle):

oh oops, supposed to be 2theta + sin 2 theta. ive corrected. so dx/dt would be 2 + 2cos2theta?

OpenStudy (irishboy123):

y

OpenStudy (lxelle):

?

OpenStudy (irishboy123):

yes

OpenStudy (lxelle):

Dy/dt = 2sin2theta right?

OpenStudy (lxelle):

@perl helppp

OpenStudy (trojanpoem):

Chain rule.

OpenStudy (perl):

$$ \Large \frac{dy}{dx }=\frac{\frac{dy}{d\theta}}{\frac {dx}{d\theta}}= \frac{2sin(2\theta)}{1 + 2 cos 2 \theta} $$

OpenStudy (perl):

but that doesnt equal tan theta. i think i copied your question wrong

OpenStudy (lxelle):

@perl oh it's 2sin2theta / 2 + 2 cos 2 theta

OpenStudy (perl):

how did you get 2 + 2 cos 2 theta for the denominator? x = theta + sin(2 theta)

OpenStudy (perl):

x = theta + sin(2 theta) dx/dtheta = 1 + 2 cos(2 theta) can you double check your directions , did I copy correctly

OpenStudy (lxelle):

I've edited the ques, supposedly 2theta + sin2theta.

OpenStudy (perl):

ok i see now :)

OpenStudy (perl):

i am using two identities sin (2u) = 2 sin u cos u cos(2u) = 2cos^2(u) - 1

OpenStudy (lxelle):

Uh huh

OpenStudy (perl):

$$ \Large{ \frac{dy}{dx }=\frac{\frac{dy}{d\theta}}{\frac {dx}{d\theta}}= \frac{2\sin(2\theta)}{2 + 2 \cos 2 \theta}\\= \frac{2\sin(2\theta)}{2(1 + \cos 2 \theta) }=\frac{\sin(2\theta)}{1 + \cos (2 \theta) }\\ = \frac{\sin(2\theta)}{1 + \cos(2\theta) } =\frac{2\sin(\theta)\cos(\theta)}{1 + (2\cos^2\theta-1) } =\frac{2\sin(\theta)\cos(\theta)}{2\cos^2\theta }\\ =\frac{\sin(\theta)}{\cos (\theta) }= \tan \theta \\ \\ } $$

OpenStudy (lxelle):

Thanksss